OpenStudy (chrisylove):

what is 20-8(q-11)=-12

OpenStudy (seratul):

First, you have to multiply the -8 by q-11. So first do -8 times q and -8 times -11.

OpenStudy (chrisylove):

20+8q-88=-12

OpenStudy (seratul):

No, it would be 20-8q=8=11 because you have to multiply -8 by q and -8 by -11.

OpenStudy (danjs):

watch the signs 20 - 8*q - 8*-11 = -12 20 - 8q + 88 = -12

OpenStudy (seratul):

I mean 20-8q+88=-12

OpenStudy (chrisylove):

where did you get the negative from

OpenStudy (seratul):

Since it is 20 - 8 (q-11), the negative is part of the 8.

OpenStudy (danjs):

20 + (-8)*(q - 11) = -12

OpenStudy (chrisylove):

im confused

OpenStudy (chrisylove):

im confused

OpenStudy (mathmale):

Your original post begins with "What is ... " These instructions are incomplete. i think you meant, "What is the value of q if .... " Please verify this by checking the source of this problem.

OpenStudy (mathmale):

Please ask specific questions, rather than saying, "I'm confused."

OpenStudy (maxine):

I created a step-by-step solution for you at Calorful: http://goo.gl/dGD3Hc This is a step by step calculator for algebra questions.

OpenStudy (alma):

I created a step-by-step solution for you at https://www.calorful.com/algebra/?q=20-8(q-11)%3D-12 This is a step by step calculator for algebra questions. “

OpenStudy (djmaster):

I created a step-by-step solution for you at Calorful: http://goo.gl/zTRDs9 This is a step by step calculator for algebra questions.